a+b=10.6
a+c=8.5
b+c=8.1
how much are a, b, c worth?
Please help .8th grade

Answer by Earlsdon(6103)

You can solve this sytem of equations by the substution method.
1) a + b = 10.6 or b = 10.6 - a
2) a + c = 8.5 or c = 8.5 - a
3) b + c = 8.1 Substitute the two above equations for b and c, then solve for a.
(10.6 - a) + (8.5 - a) = 8.1 Simplify and solve for a.
10.6 + 8.5 - 2a = 8.1 Add 2a to both sides.
19.1 = 8.1 + 2a Subtract 8.1 from both sides.
2a = 11 Divide both sides by 2
a = 5.5
b = 10.6 - a
b = 10.6 - 5.5
b = 5.1
c = 8.5 - a
c = 8.5 - 5.5
c = 3
a = 5.5
b = 5.1
c = 3